mikewwe13:

Which probability distribution table corresponds with this frequency distribution table?

mikewwe13:

mikewwe13:

mikewwe13:

mikewwe13:

mikewwe13:

loveboy:

What do you think?

mikewwe13:

probably a or C .. I'm not sure

loveboy:

Let's figure out for x = 1

loveboy:

Total freq = 11+1+2+4+2 = 20

loveboy:

Prob(x=1) = 11/20 = 0.55

loveboy:

which makes Option C correct one :)